Prev Next

SW&IT Meeting Minutes 2021-09-20

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210906
LR20210906 - Bec, Matthieu 72.0
LR20210906 - Mechtley, Brandon 80.0
LR20210906 - Eychaner, Glenn 72.0
LR20210906 - Edmeades, David 48.0
LR20210906 - Grenz, Paul 19.0
LR20210906 - Hooper, Stephen 64.0
LR20210906 - Zhang, Yang 72.0
LR20210906 - Zhang, Xianyu 96.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • luci.luci md125?
  • new observer: Ester Marini - wiki, others?

Team reports

MB

  • Last week
    • OC, Management
    • IT8470 - follow-ups, between JHill and Ilya TBD GCS improvements for saturated images
    • dnf update some dms packages that broke the OSA scripts
    • dms m1 surf - c00 and total correction views (some imperfections left for later)
    • investigate 'top' RAM reports, eventually lead nowhere
    • AO SX investigations
      • dug out AO usb drive data from DMiller (thank SH), other CBs saved by Xianyu (thank BM) - nothing there.
      • Reading more exisiting docs and reports.
      • posted comments from data we already have and presented, and some other unanswered questions
      • submitted a couple of github issues (logs, filotto), assigned to BM
    • chat with GB about http://info, bi-regular as normal
  • This week
    • 1:1 Adriano
    • ATT PM 1:many
    • git/github "policy" (= philosophy slides) for new staff
    • discuss moving forward with ao-analysis moving to Tucson, centos7 to fedora, AO data flow and retention jobs, etc. with BM and IT
    • SWG reports, overdue

GCE

  • Last week
    • Tested high priority LUCI issues with Jenny. Found that most of them appear to be problems with RTD Aladin instance and/or threading (event thread lockup).
      • The observed problem with the automatic slit alignment (IT 7642) was confirmed and may require on-sky time to fix.
    • Began investigating LUCI threading and Aladin issues.
    • Thoroughly re-read LUCI handover documents in Google Drive, now that I have a better understanding of the instrument.
      • I now understand what would be needed to test running LUCI from lucix; this is a topic for either a LUCI downtime (as discussed during the All-Hands) or 2022 Summer.
    • Set aside the ahs-luci integration and IT 8454 for the time being to work on the LUCI code fixes, which have a greater observing impact.
  • This week
    • Begin fixing and testing LUCI threading and RTD Aladin instance.

DE

  • Post-vacation catchup
  • Jim's issue down to bad dock?
  • Ester Marini stuff
  • Got all parts for backup server upgrade, will find a time to come in
  • MS Server should be sorted out; waiting for receipt of license
  • Word on EOY expenditures?

YZ

  • Previous:
    • Some MODS stuff discussion and fix attempt (modsIDL).
    • Tested Python 2.7 inotify programs (should work okay, but a few things to discuss).
    • Start to work on a MODS WCS program.
    • Initialized the new mods-obs repository (more tweaks to follow).
    • TMS meeting and some TMS data mining and analaysis.
    • Worked with Paul on Argos swingarm test in TCS/OSS.
  • Next:
    • Continue some misc LBC/MODS things (modsAlign/IDL/WCS/new github repo etc).
    • Discuss a few issues about the Python 2.7 inotify program and then move forward to develop it.
    • More TMS data discussion and some software work.
    • Work with Paul on Argos swingarm test if we can get some telescope time.

BM

  • Last week:
    • Added tool for getting data from AdSec diagnostic dumps to ao-datasystem.
    • OC investigation "data support" (cross referencing elevation, diagnostic dumps)
    • Continued getting familiar with ao-supervisor code, how and when diagnostic dumps are produced, deleted, backed up. Consulted Xianyu re: deleting of old diagnostic dumps.
    • AdSec gain=0 testing planning / AOG meetings
  • This week:
    • T-Th AdSec gain=0 tests.
      • Get familiar with Python/IDL communication to AdSec w/ICE.
    • Propose fixes to ao-supervisor logging issues (set up dev VM.)
    • AO diagnostic data migration to Tucson NAS process.
    • Get familiar with DMS code.

SH

  • VPN issues
  • Mountain network
  • robs stuff
  • Jim's laptop woes
    • Windows 10 reinstallation
  • UITS switches
  • iLocater pier stuff.

next:

  • Office today
    • Jim's laptop
    • Find room numbers for UITS switches
  • LDAP
  • Basecamp camera

PSG

  • Previous:
    • Added ARGOS project files to the repository.
    • Tested ARGOS swing arm code.
    • Found adjacent bug with white light source, unhandled exception when device is off causes OSS to crash.
    • Sending commands to D425 device causes the device to reset connection (terminal program works fine).
    • Selection of ARGOS swing arm on GUI won't "stick". This may be a deeper problem.
  • Next:
    • Fix errors that have been found.
    • Try to get more time on telescope to continue debugging.
    • Vacation day Friday.
Topic revision: r14 - 20 Sep 2021, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback